# Reality conditions for the KdV equation and quasi-periodic solutions in finite phase spaces

### v2

@@ 6.3. Real-valued Weierstrass-$\wp$-functions
Finally, we recall that functions $\wp_{1,1,2i-1}$ vanish at half-periods, that is at all $\Omega(I)$ and $\Omega^{\prime}(I^{\prime})$. Proposition 5 shows that all $\Omega(I)$ are real, and all $\Omega^{\prime}(I^{\prime})$ are purely imaginary, if all branch points of the spectral curve are real. According to Proposition 3, on $\mathfrak{U}_{\Re}(I^{\prime})=\Omega^{\prime}(I^{\prime})+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ functions $\wp_{1,2i-1}$ and $\wp_{1,1,2i-1}$, $i=1,\ldots$, $g$, are real-valued. According to Proposition 4, on $\mathfrak{U}_{\Im}(I)=\Omega(I)+\operatorname{\mathrm{Im}}\bar{\operatorname{\mathfrak{J}}}$ functions $\wp_{1,2i-1}$, $i=1,\ldots$, $g$, are real-valued, and $\wp_{1,1,2i-1}$, $i=1,\ldots$, $g$, attain purely imaginary values.
@@ 6.4. Summary
In $\bar{\operatorname{\mathfrak{J}}}=\operatorname{\mathbb{C}}^{g}$, where the Jacobian variety of the spectral curve is embedded, among $2^{g}$ affine subspaces parallel to the real axes subspace $\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ and obtained by half-period translations, there exists one subspace where Weierstrass-$\wp$-functions are bounded. With the choice of cycles as on fig. 1, this subspace is $\sum_{k=1}^{g}\tfrac{1}{2}\omega^{\prime}_{k}+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$. And among $2^{g}$ affine subspaces parallel to the imaginary axes subspace $\operatorname{\mathrm{Im}}\bar{\operatorname{\mathfrak{J}}}$ and obtained by half-period translations, there exists one subspace $\sum_{i=0}^{[(g-1)/2]}\tfrac{1}{2}\omega_{g-2i}+\operatorname{\mathrm{Im}}\bar{\operatorname{\mathfrak{J}}}$, where Weierstrass-$\wp$-functions are bounded. Then, by means of the addition law on the spectral curve, the domain of real-valued $\wp_{1,1}$-function is found. If all finite branch points of the spectral curve are real, as required for quasi-periodic solutions of the KdV equation, then $\sum_{k=1}^{g}\tfrac{1}{2}\omega^{\prime}_{k}+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ serves as the domain of the bounded real-valued $\wp_{1,1}$-function.
@@ 7. Non-linear waves
In this section we present effective numerical computation of quasi-periodic finite-gap solutions of the KdV equations.
@@ 7.1. Numerical computation
(76) $\frac{1}{\mathrm{b}^{2}}f(z,w)\equiv-\Big{(}\frac{y}{\mathrm{b}}\Big{)}^{2}+z^{2N+1}+\frac{c_{2N}}{\mathrm{b}^{2}}z^{2N}+\cdots+\frac{c_{N}}{\mathrm{b}^{2}}z^{N}\\ +\frac{h_{N-1}}{\mathrm{b}^{2}}z^{N-1}+\dots+\frac{h_{1}}{\mathrm{b}^{2}}z+\frac{h_{0}}{\mathrm{b}^{2}}=0.$
So $\lambda_{4g+2-2n}=h_{n}/\mathrm{b}^{2}$, and $y\mapsto y/\mathrm{b}$. The basis differentials of the first and second kind (53a), (53b) correspond to this canonical form. In this setup the sigma function is defined uniquely from heat equations, and can be obtained in the form of a series in coordinates $u$ of the Jacobian variety, and parameters $\lambda_{2i+2}$ of the curve. However, such a series converges slowly, which is not acceptable for computational purposes. The situation with Weierstrass-$\wp$-functions is even worse.
@@
(77) $\begin{split}&\wp_{i,j}(u)=\varkappa_{i,j}-\frac{\partial^{2}}{\partial u_{i}\partial u_{j}}\theta[K](\omega^{-1}u;\omega^{-1}\omega^{\prime}),\\ &\wp_{i,j,k}(u)=-\frac{\partial^{3}}{\partial u_{i}\partial u_{j}\partial u_{k}}\theta[K](\omega^{-1}u;\omega^{-1}\omega^{\prime}).\end{split}$
@@
$\begin{split}&\omega_{k}=2\int_{e_{2k-1}}^{e_{2k}}\mathrm{d}u,\qquad\quad\eta_{k}=2\int_{e_{2k-1}}^{e_{2k}}\mathrm{d}r,\\ &\omega^{\prime}_{k}=-2\sum_{i=1}^{k}\int_{e_{2i-2}}^{e_{2i-1}}\mathrm{d}u=2\sum_{i=k}^{g}\int_{e_{2i}}^{e_{2i+1}}\mathrm{d}u.\end{split}$
@@ 7.2. Genus $1$
According to (5.1), uniformization of (78) is given by
(79) $z=\wp_{1,1}(u),\qquad y=-\tfrac{1}{2}\mathrm{b}\wp_{1,1,1}(u),$
@@
Applyibg the reality conditions to (61), we find bounded real-valued solutions of (24), and so the stationary KdV equation ($N=1$):
(80) $\beta(x)=-\mathrm{b}\wp_{1,1}(-\mathrm{b}x+\tfrac{1}{2}\omega^{\prime}).$
@@
2a. Fundamental domains.             2b. Cnoidal waves.                                  Refer to caption
    Refer to caption Refer to caption
Note, if $\mathrm{b}$ is positive, than $\beta$ has maxima $-\mathrm{b}e_{1}$ at $x=n\omega/\mathrm{b}$, $n\in\operatorname{\mathbb{Z}}$, and minima $-\mathrm{b}e_{2}$ at $x=(n+\frac{1}{2})\omega/\mathrm{b}$, $n\in\operatorname{\mathbb{Z}}$. It follows from the fact that $\wp(\tfrac{1}{2}\omega^{\prime})=e_{1}$, $\wp(\tfrac{1}{2}\omega+\tfrac{1}{2}\omega^{\prime})=e_{2}$.
In terms of the Jacobi elliptic functions, we obtain the cnoidal wave, introduced by Korteweg and de Vries in [16]:
$-\tfrac{1}{2}\wp(-\tfrac{1}{2}x+\tfrac{1}{2}\omega^{\prime})$ $+\tfrac{1}{6}(e_{1}+e_{2}+e_{3})$
$=-\tfrac{1}{2}\bigg{(}e_{2}+(e_{3}-e_{1})\frac{\operatorname{dn}\big{(}\imath K^{\prime}\mp\tfrac{1}{2}x\sqrt{e_{3}-e_{1}};k\big{)}^{2}}{\operatorname{sn}\big{(}\imath K^{\prime}\mp\tfrac{1}{2}x\sqrt{e_{3}-e_{1}};k\big{)}^{2}}\bigg{)}$
$=-\tfrac{1}{2}\Big{(}e_{2}-(e_{2}-e_{1})\operatorname{cn}\big{(}\tfrac{1}{2}x\sqrt{e_{3}-e_{1}};k\big{)}^{2}\Big{)},$
@@ 7.3. Genus 2
(81) $-y^{2}+\mathrm{b}^{2}z^{5}+c_{3}z^{3}+c_{2}z^{2}+h_{1}z+h_{0}=0.$
Requiring real-valued solutions, we suppose that all parameters $\mathrm{b}$, $c_{3}$, $c_{2}$, $h_{1}$, $h_{0}$ are real, and the spectral curve has five real branch points $e_{1}<e_{2}<e_{3}<e_{4}<e_{5}$. According to (5.1), uniformization of (81) is given by
(82) $z^{2}=z\wp_{1,1}(u)+\wp_{1,3}(u),\qquad y=-\tfrac{1}{2}\mathrm{b}\big{(}z\wp_{1,1,1}(u)+\wp_{1,1,3}(u)\big{)},$
@@
(83) $\beta(x)=-\mathrm{b}\wp_{1,1}(-\mathrm{b}x+C_{1},-\mathrm{b}t+C_{3}),$
@@
where $(\omega^{\prime}_{1,k},\omega^{\prime}_{3,k})^{t}=\omega^{\prime}_{k}$. First indices of the entries $\omega^{\prime}_{2i-1,k}$ correspond to the labels of holomorphic differentials $\mathrm{d}u_{2i-1}$, $i=1$, $2$, and $\omega^{\prime}_{1}$, $\omega^{\prime}_{2}$ are two columns of the period matrix $\omega^{\prime}$. On the subspace $-(\mathrm{b}x,\mathrm{b}t)^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}$ the function $\wp_{1,1}$ is real-valued and bounded, with the critical values:
@@
The corresponding shapes of $\beta$, according to (83) with $\mathrm{b}=1/2$, are presented on fig. 4 and 5. On fig. 4 the reader finds shapes of $\beta(x,0)$, $x\in[0,-20(\omega_{1,1}+\omega_{1,2})]$ (left column), and $\beta(0,t)$, $t\in[0,20]$ (right column). On fig. 5 shapes of $\beta(x,t)$, $x\in[0,-8(\omega_{1,1}+\omega_{1,2})]$, $t\in[0,-4\omega_{3,2}]$ are presented for the chosen values of $h_{1}$ and $h_{0}$. Dots on fig. 5 indicate critical values (84).
Figure 4. $N=2$. Quasi-periodic waves $\beta(x,0)$ (left) and $\beta(0,t)$ (right).
@@
Figure 5. $N=2$. Quasi-periodic waves $\beta(x,t)$.
@@ 7.4. Genus $3$
(85) $-y^{2}+\mathrm{b}^{2}z^{7}+c_{5}z^{5}+c_{4}z^{4}+c_{3}z^{3}+h_{2}z^{2}+h_{1}z+h_{0}=0,$
with all real parameters, chosen in such a way that all seven branch points are real. According to (5.1), uniformization of (85) is given by
(86) $\begin{split}&z^{3}=z^{2}\wp_{1,1}(u)+z\wp_{1,3}(u)+\wp_{1,5}(u),\\ &y=-\tfrac{1}{2}\mathrm{b}\big{(}z^{2}\wp_{1,1,1}(u)+z\wp_{1,1,3}(u)+\wp_{1,1,5}(u)\big{)},\end{split}$
@@
(87) $\beta(x)=-\mathrm{b}\wp_{1,1}(-\mathrm{b}x+C_{1},-\mathrm{b}t+C_{3},C_{5}),$
@@
$\begin{split}&\qquad C_{1}=\tfrac{1}{2}\omega^{\prime}_{1,1}+\tfrac{1}{2}\omega^{\prime}_{1,2}+\tfrac{1}{2}\omega^{\prime}_{1,3},\\ &\qquad C_{3}=\tfrac{1}{2}\omega^{\prime}_{3,1}+\tfrac{1}{2}\omega^{\prime}_{3,2}+\tfrac{1}{2}\omega^{\prime}_{3,3},\\ &\qquad C_{5}=-\mathrm{c}+\tfrac{1}{2}\omega^{\prime}_{5,1}+\tfrac{1}{2}\omega^{\prime}_{5,2}+\tfrac{1}{2}\omega^{\prime}_{5,3},\quad\mathrm{c}\in\operatorname{\mathbb{R}},\end{split}$
@@
However, projection to the subspace $-(\mathrm{b}x,\mathrm{b}t,\mathrm{c})^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3}$, which serves as the domain of the quasi-periodic KdV solution $\beta$, could contain not more than one of the values (88), if $\mathrm{c}$ is one of half-periods constructed from $\tfrac{1}{2}\omega_{1}$, $\tfrac{1}{2}\omega_{2}$, $\tfrac{1}{2}\omega_{3}$.
Let $\mathcal{M}_{3}^{\circ}$ be fixed by $14\mathrm{b}^{2}=$ 1, $c_{6}=0$, $c_{5}=-84$, $c_{4}=-160$, $c_{3}=7250$. Let the hamiltonians $h_{2}$, $h_{1}$, $h_{0}$ attain six values with different mutual positions of branch points. Computed by (87) with $\mathrm{b}=1/2$ and $\mathrm{c}=0$, shapes of $\beta(x,0)$, $x\in[0,-20(\omega_{1,1}+\omega_{1,3})]$, and $\beta(0,t)$, $t\in[0,10]$ or $t\in[0,20]$, are presented below.
@@
$h_{2}=21460$, $h_{1}=-120322$, $h_{0}=-287405;\qquad$ $e_{1}\approx-12.88$, $e_{2}\approx-9.31$, $e_{3}\approx-6.21$ $e_{4}\approx-2.16$, $e_{5}\approx 4.31$, $e_{6}\approx 10.58$, $e_{7}\approx 15.67$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$h_{2}=22460$, $h_{1}=-81268$, $h_{0}=-267380\qquad$ $e_{1}\approx-12.21$, $e_{2}\approx-10.94$, $e_{3}\approx-3.57$ $e_{4}\approx-3.56$, $e_{5}\approx 3.62$, $e_{6}\approx 11.23$, $e_{7}\approx 15.44$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
On fig. 6 shapes of $\beta(x,t)$, $x\in[0,6]$, $t\in[0,0.6]$ at some chosen values of $h_{2}$, $h_{1}$, $h_{0}$ are presented.
Figure 6. $N=3$. Quasi-periodic waves $\beta(x,t)$, $x\in[0,6]$, $t\in[0,0.6]$.

### v4

@@ 6.3. Real-valued Weierstrass-$\wp$-functions
Proposition 1 shows that all $\Omega(I)$ are real, and all $\Omega^{\prime}(I^{\prime})$ are purely imaginary, if all branch points of the curve are real. Therefore, according to Proposition 4, on $\mathfrak{U}_{\Re}(I^{\prime})=\Omega^{\prime}(I^{\prime})+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ functions $\wp_{1,2i-1}$ and $\wp_{1,1,2i-1}$, $i=1,\ldots$, $g$, are real-valued. According to Proposition 5, on $\mathfrak{U}_{\Im}(I)=\Omega(I)+\operatorname{\mathrm{Im}}\bar{\operatorname{\mathfrak{J}}}$ functions $\wp_{1,2i-1}$, $i=1,\ldots$, $g$, are real-valued, and $\wp_{1,1,2i-1}$, $i=1,\ldots$, $g$, take only purely imaginary values.
In the finite-gap solution (48) we choose
$\bm{C}=\Omega^{\prime}(\hat{I}^{\prime})+\mathbf{c}=\sum_{k=1}^{g}\tfrac{1}{2}\omega^{\prime}_{k}+(\mathrm{c}_{1},\mathrm{c}_{3},\mathrm{c}_{5},\dots,\mathrm{c}_{2N-1})^{t},$
with $\hat{I}^{\prime}$ defined in Proposition 2, and arbitrary real $\mathrm{c}_{2n-1}$, $n=1$, …, $N$.
Remark 6.
Note that the vector of Riemann constants $K$ is expressed in terms of $\Omega^{\prime}(\hat{I}^{\prime})$ and $\Omega(\hat{I})$ defined in Propositions 2 and 3, namely
$K=\omega^{-1}\big{(}\Omega(\hat{I})+\Omega^{\prime}(\hat{I}^{\prime})\big{)}.$
Therefore, it is convenient to assign $\bm{C}=\omega K+\mathbf{c}$.
Remark 7.
Comparing the KdV solution in the form of (48) with (2), we find the following correspondence:
$\bm{U}=-\tfrac{1}{2}\omega^{-1}_{1},\quad\bm{W}=-\tfrac{1}{2}\omega^{-1}_{2},\quad\bm{D}=\omega^{-1}\bm{C},\quad c=-\tfrac{1}{4}\varkappa_{1,1},$
where $\omega^{-1}_{k}$ denotes the $k$-th column of the matrix $\omega^{-1}$, and $\varkappa=\eta\omega^{-1}$.
@@ 6.4. Summary
In $\bar{\operatorname{\mathfrak{J}}}=\operatorname{\mathbb{C}}^{g}$, where the Jacobian variety of the spectral curve is embedded, among $2^{g}$ affine subspaces $\mathfrak{U}_{\Re}$ parallel to the real axes subspace $\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ and obtained by half-period translations, there exists one subspace where Weierstrass-$\wp$-functions are bounded. With the choice of cycles as on fig. 1, this subspace is $\sum_{k=1}^{g}\tfrac{1}{2}\omega^{\prime}_{k}+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$. And among $2^{g}$ affine subspaces $\mathfrak{U}_{\Im}$ parallel to the imaginary axes subspace $\operatorname{\mathrm{Im}}\bar{\operatorname{\mathfrak{J}}}$ and obtained by half-period translations, there exists one subspace $\sum_{i=0}^{[(g-1)/2]}\tfrac{1}{2}\omega_{g-2i}+\operatorname{\mathrm{Im}}\bar{\operatorname{\mathfrak{J}}}$, where Weierstrass-$\wp$-functions are bounded. Then, by means of the addition law on the curve, it was proven that functions $\wp_{1,2i-1}$ on all $\mathfrak{U}_{\Re}$ and $\mathfrak{U}_{\Im}$ subspaces are real-valued, and functions $\wp_{1,1,2i-1}$ are real-valued on subspaces $\mathfrak{U}_{\Re}$, and purely imaginary-valued on subspaces $\mathfrak{U}_{\Im}$. Thus, if all finite branch points of the spectral curve are real, as required for quasi-periodic solutions of the KdV equation, then $\sum_{k=1}^{g}\tfrac{1}{2}\omega^{\prime}_{k}+\operatorname{\mathrm{Re}}\bar{\operatorname{\mathfrak{J}}}$ serves as the domain of the bounded real-valued $\wp_{1,1}$-function.
@@ 7. Non-linear waves
In this section we present effective numerical computation of quasi-periodic finite-gap solutions of the KdV equation.
An analytical approach to computation of Weierstrass-$\wp$-functions is used, for more details see [8]. Once a curve is chosen, periods of the first kind (59), and the corresponding periods of the second kind are computed. Then Weierstrass-$\wp$-functions are calculated by the formulas (58). All computations and graphical representation are performed in Wolfram Mathematica 12. The function NIntegrate is used for numerical integration of periods, and NSolve for computing branch points in genera $2$ and $3$. The default precision is applied, which is sufficient for graphical representation. The theta function is approximated by a partial sum of (6), $|n_{i}|\leqslant 5$ is sufficient.
@@ 7.1. Numerical computation
(57) $0=\frac{1}{\mathrm{b}^{2}}F(z,w)\equiv-\Big{(}\frac{w}{\mathrm{b}}\Big{)}^{2}+z^{2N+1}+\frac{c_{2N}}{\mathrm{b}^{2}}z^{2N}+\cdots+\frac{c_{N}}{\mathrm{b}^{2}}z^{N}\\ +\frac{h_{N-1}}{\mathrm{b}^{2}}z^{N-1}+\dots+\frac{h_{1}}{\mathrm{b}^{2}}z+\frac{h_{0}}{\mathrm{b}^{2}},$
where $z=x$, $w/\mathrm{b}=y$, and so $\lambda_{4g+2-2n}=h_{n}/\mathrm{b}^{2}$. The basis differentials of the first and second kind (40) correspond to this canonical form. In this setup the sigma function is defined uniquely from heat equations, and can be obtained in the form of a series in coordinates $u$ of the Jacobian variety, and parameters $\lambda_{2i+2}$ of the curve. However, such a series converges slowly, which is not acceptable for computational purposes. The situation with Weierstrass-$\wp$-functions is even worse.
@@
(58) $\begin{split}&\wp_{i,j}(u)=\varkappa_{i,j}-\frac{\partial^{2}}{\partial u_{i}\partial u_{j}}\log\theta[K](\omega^{-1}u;\omega^{-1}\omega^{\prime}),\\ &\wp_{i,j,k}(u)=-\frac{\partial^{3}}{\partial u_{i}\partial u_{j}\partial u_{k}}\log\theta[K](\omega^{-1}u;\omega^{-1}\omega^{\prime}).\end{split}$
@@
(59) $\begin{split}&\omega_{k}=2\int_{e_{2k-1}}^{e_{2k}}\mathrm{d}u,\qquad\quad\eta_{k}=2\int_{e_{2k-1}}^{e_{2k}}\mathrm{d}r,\\ &\omega^{\prime}_{k}=-2\sum_{i=1}^{k}\int_{e_{2i-2}}^{e_{2i-1}}\mathrm{d}u=2\sum_{i=k}^{g}\int_{e_{2i}}^{e_{2i+1}}\mathrm{d}u.\end{split}$
@@ 7.2. Genus $1$
According to (41), uniformization of (60) is given by
(61) $z=\wp_{1,1}(u),\qquad w=-\tfrac{1}{2}\mathrm{b}\wp_{1,1,1}(u),$
@@
Applying the reality conditions to (48), we find bounded real-valued solutions of (28), and so the stationary KdV equation ($N=1$):
(62) $\beta(\mathrm{x})=-\mathrm{b}\wp_{1,1}(-\mathrm{b}\mathrm{x}+\tfrac{1}{2}\omega^{\prime}).$
@@
2a. Fundamental domains.             2b. Cnoidal waves.                                  Refer to caption     Refer to caption Refer to caption
Note, if $\mathrm{b}$ is positive, than $\beta$ has maxima $-\mathrm{b}e_{1}$ at $\mathrm{x}=n\omega/\mathrm{b}$, $n\in\operatorname{\mathbb{Z}}$, and minima $-\mathrm{b}e_{2}$ at $\mathrm{x}=(n+\frac{1}{2})\omega/\mathrm{b}$, $n\in\operatorname{\mathbb{Z}}$. It follows from the fact that $\wp(\tfrac{1}{2}\omega^{\prime})=e_{1}$, $\wp(\tfrac{1}{2}\omega+\tfrac{1}{2}\omega^{\prime})=e_{2}$.
In terms of the Jacobi elliptic functions, we obtain the cnoidal wave, found by Korteweg and de Vries [25]:
$-\tfrac{1}{2}\wp(-\tfrac{1}{2}\mathrm{x}+\tfrac{1}{2}\omega^{\prime})$ $+\tfrac{1}{6}(e_{1}+e_{2}+e_{3})$
$=-\tfrac{1}{2}\bigg{(}e_{2}+(e_{3}-e_{1})\frac{\operatorname{dn}\big{(}\imath K^{\prime}\mp\tfrac{1}{2}\mathrm{x}\sqrt{e_{3}-e_{1}};k\big{)}^{2}}{\operatorname{sn}\big{(}\imath K^{\prime}\mp\tfrac{1}{2}\mathrm{x}\sqrt{e_{3}-e_{1}};k\big{)}^{2}}\bigg{)}$
$=-\tfrac{1}{2}\Big{(}e_{2}-(e_{2}-e_{1})\operatorname{cn}\big{(}\tfrac{1}{2}\mathrm{x}\sqrt{e_{3}-e_{1}};k\big{)}^{2}\Big{)},$
@@ 7.3. Genus 2
(63) $-w^{2}+\mathrm{b}^{2}z^{5}+c_{3}z^{3}+c_{2}z^{2}+h_{1}z+h_{0}=0.$
Requiring real-valued solutions, we suppose that all parameters $\mathrm{b}$, $c_{3}$, $c_{2}$, $h_{1}$, $h_{0}$ are real, and the spectral curve has five real branch points $e_{1}<e_{2}<e_{3}<e_{4}<e_{5}$. According to (41), uniformization of (63) is given by
(64) $z^{2}=z\wp_{1,1}(u)+\wp_{1,3}(u),\qquad w=-\tfrac{1}{2}\mathrm{b}\big{(}z\wp_{1,1,1}(u)+\wp_{1,1,3}(u)\big{)},$
@@
(65) $\beta(\mathrm{x},\mathrm{t})=-\mathrm{b}\wp_{1,1}(-\mathrm{b}\mathrm{x}+C_{1},-\mathrm{b}\mathrm{t}+C_{3}),$
@@
where $(\omega^{\prime}_{1,k},\omega^{\prime}_{3,k})^{t}=\omega^{\prime}_{k}$. First indices of the entries $\omega^{\prime}_{2i-1,k}$ correspond to the labels of holomorphic differentials $\mathrm{d}u_{2i-1}$, $i=1$, $2$, and $\omega^{\prime}_{1}$, $\omega^{\prime}_{2}$ are two columns of the period matrix $\omega^{\prime}$. On the subspace $-(\mathrm{b}\mathrm{x},\mathrm{b}\mathrm{t})^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}$ the function $\wp_{1,1}$ is real-valued and bounded, with the critical values:
@@
The corresponding shapes of $\beta$, according to (65) with $\mathrm{b}=1/2$, are presented on fig. 4 and 5. On fig. 4 the reader finds shapes of $\beta(\mathrm{x},0)$, $\mathrm{x}\in[0,-20(\omega_{1,1}+\omega_{1,2})]$ (left column), and $\beta(0,\mathrm{t})$, $\mathrm{t}\in[0,20]$ (right column). On fig. 5 shapes of $\beta(\mathrm{x},\mathrm{t})$, $\mathrm{x}\in[0,-8(\omega_{1,1}+\omega_{1,2})]$, $\mathrm{t}\in[0,-4\omega_{3,2}]$ are presented for the chosen values of $h_{1}$ and $h_{0}$. Dots on fig. 5 indicate critical values (66).
Figure 4. $N=2$. Quasi-periodic waves $\beta(\mathrm{x},0)$ (left) and $\beta(0,\mathrm{t})$ (right).
@@
Figure 5. $N=2$. Quasi-periodic waves $\beta(\mathrm{x},\mathrm{t})$.
@@ 7.4. Genus $3$
(67) $-w^{2}+\mathrm{b}^{2}z^{7}+c_{5}z^{5}+c_{4}z^{4}+c_{3}z^{3}+h_{2}z^{2}+h_{1}z+h_{0}=0,$
with all real parameters, chosen in such a way that all seven branch points are real. According to (41), uniformization of (67) is given by
(68) $\begin{split}&z^{3}=z^{2}\wp_{1,1}(u)+z\wp_{1,3}(u)+\wp_{1,5}(u),\\ &w=-\tfrac{1}{2}\mathrm{b}\big{(}z^{2}\wp_{1,1,1}(u)+z\wp_{1,1,3}(u)+\wp_{1,1,5}(u)\big{)},\end{split}$
@@
(69) $\beta(\mathrm{x},\mathrm{t})=-\mathrm{b}\wp_{1,1}(-\mathrm{b}\mathrm{x}+C_{1},-\mathrm{b}\mathrm{t}+C_{3},C_{5}),$
@@
$\begin{split}&\qquad C_{1}=\tfrac{1}{2}\omega^{\prime}_{1,1}+\tfrac{1}{2}\omega^{\prime}_{1,2}+\tfrac{1}{2}\omega^{\prime}_{1,3},\\ &\qquad C_{3}=\tfrac{1}{2}\omega^{\prime}_{3,1}+\tfrac{1}{2}\omega^{\prime}_{3,2}+\tfrac{1}{2}\omega^{\prime}_{3,3},\\ &\qquad C_{5}=-\mathrm{c}_{5}+\tfrac{1}{2}\omega^{\prime}_{5,1}+\tfrac{1}{2}\omega^{\prime}_{5,2}+\tfrac{1}{2}\omega^{\prime}_{5,3},\quad\mathrm{c}\in\operatorname{\mathbb{R}},\end{split}$
@@
However, projection to the subspace $-(\mathrm{b}\mathrm{x},\mathrm{b}\mathrm{t},\mathrm{c})^{t}+\tfrac{1}{2}\omega^{\prime}_{1}+\tfrac{1}{2}\omega^{\prime}_{2}+\tfrac{1}{2}\omega^{\prime}_{3}$, which serves as the domain of the quasi-periodic KdV solution $\beta$, could contain not more than one of the values (70), if $\mathrm{c}$ is one of half-periods constructed from $\tfrac{1}{2}\omega_{1}$, $\tfrac{1}{2}\omega_{2}$, $\tfrac{1}{2}\omega_{3}$.
Let $\mathcal{M}_{3}^{\circ}$ be fixed by $14\mathrm{b}^{2}=$ 1, $c_{6}=0$, $c_{5}=-84$, $c_{4}=-160$, $c_{3}=7250$. Let the hamiltonians $h_{2}$, $h_{1}$, $h_{0}$ attain six values with different mutual positions of branch points. Computed by (69) with $\mathrm{b}=1/2$ and $\mathrm{c}_{5}=0$, shapes of $\beta(\mathrm{x},0)$, $\mathrm{x}\in[0,-20(\omega_{1,1}+\omega_{1,3})]$, and $\beta(0,\mathrm{t})$, $\mathrm{t}\in[0,10]$ or $\mathrm{t}\in[0,20]$, are presented below.
@@
$h_{2}=21460$, $h_{1}=-120322$, $h_{0}=-287405;\qquad$ $e_{1}\approx-12.88$, $e_{2}\approx-9.31$, $e_{3}\approx-6.21$ $e_{4}\approx-2.16$, $e_{5}\approx 4.31$, $e_{6}\approx 10.58$, $e_{7}\approx 15.67$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $h_{2}=22460$, $h_{1}=-81268$, $h_{0}=-267380\qquad$ $e_{1}\approx-12.21$, $e_{2}\approx-10.94$, $e_{3}\approx-3.57$ $e_{4}\approx-3.56$, $e_{5}\approx 3.62$, $e_{6}\approx 11.23$, $e_{7}\approx 15.44$;     [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
On fig. 6 shapes of $\beta(\mathrm{x},\mathrm{t})$, $\mathrm{x}\in[0,6]$, $\mathrm{t}\in[0,0.6]$ at some chosen values of $h_{2}$, $h_{1}$, $h_{0}$ are presented.
Figure 6. $N=3$. Quasi-periodic waves $\beta(\mathrm{x},\mathrm{t})$, $\mathrm{x}\in[0,6]$, $\mathrm{t}\in[0,0.6]$.
[…]
$\phantom{mm}h_{0}=111126$.
[…]
$\phantom{mm}h_{0}=-338529$.
6c. $h_{2}=12678$, $h_{1}=-172935$,
$\phantom{mm}h_{0}=-12782$.
6c. $h_{2}=21460$, $h_{1}=-120322$
$\phantom{mm}h_{0}=-287405$.
6e. $h_{2}=22460$, $h_{1}=-81268$,
$\phantom{mm}h_{0}=-267380$.
6f. $h_{2}=32678$, $h_{1}=-63757$,
$\phantom{mm}h_{0}=-372539$.